# Low-Rank Key Value Attention

### v1

@@ Abstract
Transformer pretraining is increasingly constrained by memory and compute requirements, with the key–value (KV) cache emerging as a dominant bottleneck during training and autoregressive decoding. We propose low-rank KV adaptation (LRKV), a simple modification of multi-head attention that reduces KV cache memory by exploiting redundancy across attention heads while preserving full token-level resolution. Each layer uses a shared full-rank KV projection augmented with low-rank, head-specific residuals, yielding a continuous trade-off between complete sharing and fully independent attention.
LRKV is a drop-in replacement for standard multi-head attention and directly subsumes query-sharing approaches such as multi-query and grouped-query attention, while remaining distinct from latent-compression methods such as multi-latent attention (MLA). Across large-scale pretraining experiments, LRKV consistently achieves faster loss reduction, lower validation perplexity, and stronger downstream task performance than standard attention, MQA/GQA, and MLA. At the 2.5B scale, LRKV outperforms standard attention while using roughly half the KV cache, and reaches equivalent model quality with up to 20–25% less training compute when measured in cumulative FLOPs. To explain these gains, we analyze attention head structure in operator space and show that LRKV preserves nearly all functional head diversity relative to standard attention, whereas more aggressive KV-sharing mechanisms rely on compensatory query specialization. Together, these results establish LRKV as a practical and effective attention mechanism for scaling Transformer pretraining under memory- and compute-constrained regimes.
@@ 1 Introduction
Transformers are the dominant architecture for large-scale sequence modeling in language, vision, and multimodal domains (Vaswani et al., 2017; OpenAI, 2023), but as their size, sequence length, and context window grow, so does, rapidly, their computational and memory costs. KV-caching, which stores the atttention key and value representations, is a primary contributor to this overhead, as it spans every attention layer, and scales linearly with sequence length and count of heads. The cumulative KV footprint of modern models with tens of billion parameters can exceed the parameter memory itself, especially for long-context inference (Dao et al., 2024; Child et al., 2019).
A range of approaches have been proposed to alleviate the growing KV cache cost. Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped-Query Attention (GQA) (Ainslie et al., 2023) reduce memory and latency by sharing key-value (KV) projections across heads or groups of heads, and are now standard in large-scale models such as PaLM and LLaMA (Touvron et al., 2023; 2024). However, aggressive KV sharing introduces a fundamental trade-off: while memory is reduced, head-level representational diversity is constrained, even though distinct attention heads are known to encode complementary syntactic and semantic patterns (Clark et al., 2019; Michel et al., 2019).
@@
Figure 1: Comparison of attention mechanisms and their KV cache costs. (a) Multi-Head Attention (MHA) uses independent K/V projections per head with cache cost $2LHd_{h}$. (b) Multi-Latent Attention compresses inputs to latent dimension $d_{c}$ before per-head projections (cache: $Ld_{c}$ where $d_{c}\ll d$). (c) Multi-Query Attention (MQA) shares a single K/V across all heads (cache: $2Ld_{h}$) but suffers from low diversity. (d) Grouped-Query Attention (GQA) shares K/V within groups of heads (cache: $2L(H/G)d_{h}$), balancing efficiency and diversity. (e) Our proposed LR-KV method maintains shared full-rank $W_{K},W_{V}$ projections while each head adds a low-rank residual $U_{h}(B_{h})^{\top}$ (rank $r\ll d_{h}$). We cache both the shared features and per-head latents $R_{h}=XU_{h}$, achieving cache cost $2L(d_{h}+Hr)$. Dashed boxes indicate cached components. LR-KV provides head diversity while maintaining cache efficiency comparable to MLA.
@@
Beyond KV sharing, efficiency research has pursued complementary directions such as sparse or kernelized attention(Wang et al., 2020; Beltagy et al., 2020; Child et al., 2019), architectural optimizations(Dao et al., 2024) , and latent compression methods including Multi-Latent Attention (MLA)(Liu et al., 2024a). However, none of these approaches explicitly resolve the duplication of per-head key and value representations that dominates the KV memory footprint in large models.
@@
In this work, we propose Low-Rank KV Adaptation (LRKV), a simple and effective modification of multi-head attention that directly addresses this tension by exploiting structured redundancy across heads while preserving head specialization. Each Transformer layer maintains a single shared, full-rank key and value projection serving as a global basis across heads, while each head learns a compact, trainable low-rank residual that perturbs the shared representations. The shared component encodes global relational structure, and the low-rank residuals restore the localized specialization that independent heads would otherwise provide. This factorization significantly reduces KV memory while retaining the head-level diversity critical for expressivity.

### v2

@@ Abstract
The key-value (KV) cache is a primary memory bottleneck in Transformers. We propose Low-Rank Key-Value (LRKV) attention, which reduces KV cache memory by exploiting redundancy across attention heads, while being compute efficient. Each layer uses a shared full-rank KV projection augmented with low-rank, head-specific residuals, providing a continuous trade-off between complete sharing and full independence. After pretraining models of size 128M to 6.3B parameters, LRKV consistently achieves the lowest test loss among standard MHA, MQA/GQA, and MLA while using only 45-53% of MHA’s KV cache. LRKV reaches equivalent baseline quality 18-25% faster (measured in training steps). After supervised midtraining, LRKV achieves the highest downstream task performance across ARC-Easy, ARC-Challenge, MMLU, GSM8K, and HumanEval benchmarks.
Machine Learning, ICML
@@ 1 Introduction
[…]
Figure 1: Comparison of attention mechanisms. Standard MHA: Uses $H$ independent projections per head, requiring high cache cost ($2LHd_{h}$). (MQA/GQA: Shares projections across all heads, reducing cache but losing head-specific detail. LRKV (Ours): Combines shared full-rank parameters with head-specific low-rank residuals ($\mathbf{U}_{h}\mathbf{B}_{h}^{T}$, rank $r\ll d_{h}$), achieving low cache cost $2L(d_{h}+Hr)$ while preserving head diversity.
A range of approaches have been proposed to alleviate the growing KV cache cost. Multi-Query Attention (Shazeer, 2019, MQA) and Grouped-Query Attention (Ainslie et al., 2023, GQA) reduce memory and latency by sharing key-value (KV) projections across heads or groups of heads, and are now standard in large-scale models such as PaLM and LLaMA (Touvron et al., 2023, 2024). However, aggressive KV sharing introduces a fundamental trade-off: while memory is reduced, head-level representational diversity is constrained, even though distinct attention heads are known to encode complementary syntactic and semantic patterns (Clark et al., 2019; Michel et al., 2019).
@@
Beyond KV sharing, efficiency research has pursued complementary directions such as sparse or kernelized attention (Wang et al., 2020; Beltagy et al., 2020; Child et al., 2019), architectural optimizations (Dao et al., 2024) , and latent compression methods including Multi-Latent Attention (Liu et al., 2024, MLA). However, none of these approaches explicitly resolve the duplication of per-head key and value representations that dominates the KV memory footprint in large models.
@@
In this work, we propose Low-Rank Key-Value Attention (LRKV), a simple modification of multi-head attention that exploits structured redundancy across heads while preserving head specialization. Each Transformer layer maintains a shared, full-rank key and value projection that serves as a global basis, while each head learns a compact, trainable low-rank residual that captures head-specific deviations. The shared component encodes global relational structure, and the residuals restore the localized diversity otherwise lost under aggressive KV sharing. This factorization substantially reduces KV memory while retaining the expressivity of multi-head attention.
LRKV differs fundamentally from prior KV-efficient designs in where compression is applied. Unlike LoRA (Hu et al., 2022), which introduces low-rank structure post hoc during fine-tuning, LRKV learns head-specific low-rank residuals jointly during pretraining. Unlike factorized QKV methods that reduce rank along the token or hidden dimension (Wang et al., 2020; Saxena et al., 2024; Chang et al., 2025), LRKV preserves full token resolution and compresses only across attention heads. We now discuss the main contributions of this paper.
Contributions.(1) LRKV reduces KV cache by 45-53% while achieving lowest test loss across 128M-6.3B models, reaching baseline performance 18-25% faster in training steps. (2) LRKV achieves highest downstream performance across ARC, MMLU, GSM8K, demonstrating better pretraining translates to improved capabilities. (3) Gauge-invariant analysis shows LRKV effectively uses low-rank structure by preserving 93.5% head diversity vs 94% MHA. [^1] Code and trained models will be released upon acceptance.
2 Related Work
KV-Sharing Mechanisms. MQA and GQA reduce KV cache by sharing projections across heads, but sacrifice head-level expressivity. MQA (Shazeer, 2019) uses a single shared KV projection for all heads, achieving maximal cache reduction but constraining diversity. GQA (Ainslie et al., 2023) groups heads to share KV projections within groups, interpolating between MQA and full MHA. While both methods reduce memory, they fundamentally limit the representational capacity available to each head. LRKV preserves shared projection efficiency while restoring diversity through structured low-rank head residuals, achieving better quality at comparable cache sizes.
Latent Compression Approaches. MLA (Liu et al., 2024) takes a complementary approach: rather than sharing across heads, it compresses across the token dimension by projecting inputs into a low-dimensional latent space ($d_{c}\ll d$) before caching. While effective for extreme compression, this bottleneck constrains all heads to operate through the same latent representation and requires T-dependent reconstruction overhead during generation (see  Appendix F). LRKV instead preserves full token-level resolution and compresses along the head dimension via additive factorization ($\mathbf{W}_{h}=\mathbf{W}_{\text{shared}}+\mathbf{U}_{h}\mathbf{B}_{h}^{\top}$), avoiding latent bottlenecks while maintaining per-head specialization. These represent orthogonal design choices: MLA optimizes for minimal cache size via aggressive token compression, while LRKV balances memory efficiency with head-level expressivity.
Low-Rank Parameterization. LoRA (Hu et al., 2022) introduces low-rank updates for parameter-efficient fine-tuning, intervening on the optimization pathway rather than the cached representations. LRKV applies the low-rank principle directly to KV projections during pretraining, shaping the structure of cached features from the start. Recent work explores low-rank QKV factorizations (Xie and others, 2023; Khalaf et al., 2025; Lv et al., 2024) to reduce parameters or computation; LRKV differs by preserving a full-rank shared base (capturing global structure) while using low-rank residuals only for head-specific deviations. Concurrent work on factorized KV mechanisms (MFA/MFA-KR (Hu et al., 2025), TPA (Zhang et al., 2025)) explores related structured compression ideas.
Complementary Efficiency Methods. LRKV is orthogonal to other efficiency techniques: sparse/kernelized attention (Child et al., 2019; Beltagy et al., 2020; Wang et al., 2020) reduce computational complexity but not KV cache size; architectural optimizations like FlashAttention (Dao et al., 2024) improve throughput through better memory access patterns; quantization and pruning reduce precision or parameters. For head diversity analysis, we extend prior metrics (Michel et al., 2019; Kornblith et al., 2019; Wang and Wang, 2025) using gauge-invariant bilinear forms with centered Gram matrix analysis (see Appendix A).
3 Methodology
Transformers represent each token in a sequence through queries, keys, and values that interact via scaled dot-product attention. For an input matrix $\mathbf{X}\in\mathbb{R}^{L\times d}$, the $h$-th attention head computes $\mathbf{Q}_{h}=\mathbf{X}\mathbf{W}_{h}^{Q}$, $\mathbf{K}_{h}=\mathbf{X}\mathbf{W}_{h}^{K}$, and $\mathbf{V}_{h}=\mathbf{X}\mathbf{W}_{h}^{V}$, where $\mathbf{W}_{h}^{Q,K,V}\in\mathbb{R}^{d\times d_{h}}$ and $d_{h}=d/H$ for $H$ heads. The head output is
$\mathbf{O}_{h}=\mathrm{softmax}\!\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V}_{h},$ (1)
and outputs from all heads are concatenated and projected to form the layer output.
KV caching. During autoregressive decoding, the KV cache stores $\mathbf{K}_{h}$ and $\mathbf{V}_{h}$ for all previous tokens and heads, incurring per-layer memory $M_{\text{standard}}=2LHd_{h}=2Ld$. Equivalently, per token the cache stores $2Hd_{h}$ floating-point values (keys and values across $H$ heads). Over $N$ layers in bfloat16 precision, this is $2N\cdot(2LHd_{h})$ bytes of KV storage.
Low-Rank KV Attention. We parameterize each head’s key/value projection as a shared dense (unconstrained) base plus a head-specific low-rank residual, while keeping the resulting projection shape $d\times d_{h}$:
$\mathbf{W}_{h}^{K}$ $=\mathbf{W}_{\mathrm{shared}}^{K}+\mathbf{U}_{h}^{K}{\mathbf{B}_{h}^{K}}^{\top},$ (2)
$\mathbf{W}_{h}^{V}$ $=\mathbf{W}_{\mathrm{shared}}^{V}+\mathbf{U}_{h}^{V}{\mathbf{B}_{h}^{V}}^{\top},$ (3)
where $\mathbf{W}_{\mathrm{shared}}^{K,V}\in\mathbb{R}^{d\times d_{h}}$ are dense projection matrices with no rank constraint, $\mathbf{U}_{h}^{K,V}\in\mathbb{R}^{d\times r}$, $\mathbf{B}_{h}^{K,V}\in\mathbb{R}^{d_{h}\times r}$, and $r\ll d_{h}$. The effective keys and values are $\mathbf{K}_{h}=\mathbf{X}\mathbf{W}_{h}^{K}$ and $\mathbf{V}_{h}=\mathbf{X}\mathbf{W}_{h}^{V}$. When $r=0$, LRKV reduces to complete KV sharing (MQA-style) within a layer; increasing $r$ interpolates continuously toward standard MHA.
Training and initialization. All parameters in Equation 3 are optimized jointly during pretraining. Gradients from the shared and residual paths are additive, allowing the model to learn how much per-head variation is required. We initialize $\mathbf{W}_{\mathrm{shared}}^{K,V}$ using standard Kaiming initialization for attention projections. The per-head factors $\mathbf{U}_{h}^{K,V}$ and $\mathbf{B}_{h}^{K,V}$ are initialized to small random values scaled by $1/\sqrt{r}$ such that initial residuals $\mathbf{U}_{h}{\mathbf{B}_{h}}^{\top}$ have magnitude $\approx 0.1\times\|\mathbf{W}_{\mathrm{shared}}\|_{F}$, ensuring the model starts close to the shared baseline and gradually learns head specialization. During training, we materialize full $\mathbf{K}_{h}$ and $\mathbf{V}_{h}$ matrices before attention computation for simplicity; the factored form is used only during inference to realize memory savings. We apply low-rank factorization only to key and value projections, as these are cached during inference; query projections remain independent per head at full rank.
LRKV caching scheme. During decoding, the bottleneck is storing $\mathbf{K}_{h},\mathbf{V}_{h}\in\mathbb{R}^{L\times d_{h}}$ for every head. LRKV instead caches shared features once per layer,
$\mathbf{K}_{\mathrm{shared}}=\mathbf{X}\mathbf{W}_{\mathrm{shared}}^{K},\quad\mathbf{V}_{\mathrm{shared}}=\mathbf{X}\mathbf{W}_{\mathrm{shared}}^{V}\in\mathbb{R}^{L\times d_{h}},$
and compact per-head latents
$\mathbf{R}_{h}^{K}=\mathbf{X}\mathbf{U}_{h}^{K},\quad\mathbf{R}_{h}^{V}=\mathbf{X}\mathbf{U}_{h}^{V}\in\mathbb{R}^{L\times r}.$
The full per-head features implied by Equation 3 are
$\mathbf{K}_{h}$ $=\mathbf{K}_{\mathrm{shared}}+\mathbf{R}_{h}^{K}{\mathbf{B}_{h}^{K}}^{\top},$ (4)
$\mathbf{V}_{h}$ $=\mathbf{V}_{\mathrm{shared}}+\mathbf{R}_{h}^{V}{\mathbf{B}_{h}^{V}}^{\top}.$ (5)
Importantly, the matrices $\mathbf{B}_{h}^{K,V}$ are model parameters and are not cached per token.
Attention computation without explicit reconstruction. Naively materializing $\mathbf{K}_{h},\mathbf{V}_{h}$ for all cached tokens would cost $O(Lrd_{h})$ memory and compute. Instead, we exploit associativity to compute attention exactly without forming full per-head KV tensors. For a decoding step with query $\mathbf{q}_{h}\in\mathbb{R}^{d_{h}}$,
and for attention weights $\mathbf{a}_{h}\in\mathbb{R}^{1\times L}$,
Equations 6 and 7 compute the exact attention logits and outputs implied by Equation 4Equation 5, but avoid explicit reconstruction of $\mathbf{K}_{h},\mathbf{V}_{h}$. This form can be implemented inside fused attention kernels. See Table 4 for a detailed mathematical comparison to existing mechanisms.
$M_{\text{standard}}=2LHd_{h}.$
LRKV stores shared features plus per-head latents:
$M_{\text{LRKV}}$ $=\underbrace{2Ld_{h}}_{\text{shared }(K,V)}+\underbrace{2LHr}_{\text{per-head latents}}$
$=2L(d_{h}+Hr).$ (8)
Thus the memory ratio is
Per token, LRKV stores $2d_{h}+2Hr$ values versus $2Hd_{h}$ for standard attention.
Table 1: Pretraining performance across model scales. LRKV achieves competitive test loss across all model scales (128M, 1.2B, 2.5B, 6.3B) while maintaining efficient KV cache usage.
Arch. KV Cache % 128M 1.2B 2.5B 6.3B
Model KV Heads 128M 1.2B 2.5B 6.3B CE $\downarrow$ BPB $\downarrow$ CE $\downarrow$ BPB $\downarrow$ CE $\downarrow$ BPB $\downarrow$ CE $\downarrow$ BPB $\downarrow$
Standard MHA 6/12/18/32 100 100 100 100 2.903 0.878 2.530 0.765 2.389 0.723 2.319 0.701
GQA 3/4/6/2 50 33 33 6 2.918 0.883 2.536 0.767 2.397 0.725 2.354 0.712
MQA 1 17 8 6 3 2.929 0.886 2.573 0.778 2.408 0.729 2.304 0.697
MLA 17 17 17 25 2.901 0.876 2.564 0.775 2.392 0.724 2.377 0.719
LRKV 53 48 48 45 2.893 0.872 2.509 0.758 2.376 0.719 2.288 0.692
KV Heads show values for 128M/1.2B/2.5B/6.3B scales. KV Cache % indicates memory relative to Standard MHA baseline (see Appendix G for profiling details). Test metrics on held-out FineWeb-Edu (100B tokens for 128M, 50B for 1.2B/2.5B/6.3B).
Compute overhead during decoding (FLOPs). Per decoding step, standard attention has dominant per-head computational cost $O(Ld_{h})$ from query–key dot products and value aggregation. Using the distributed forms in Equation 6 and Equation 7, LRKV adds an extra $\Delta\mathrm{FLOPs}=O(Lr+rd_{h})$ per head per decoding step. For long contexts ($L\gg 1$), the dominant additional term scales as $O(Lr)$, yielding a relative overhead of approximately $r/d_{h}$ compared to standard attention.
Memory bandwidth. Modern inference is often bandwidth-bound rather than compute-bound  (Dao et al., 2024). Per decoding step, standard MHA reads $2H$ cached tensors ($\mathbf{K}_{h}$ and $\mathbf{V}_{h}$ for each head). LRKV reads two shared tensors plus $2H$ per-head latent tensors. Since the shared tensors ($L\times d_{h}$) are reused across heads and the per-head latents ($L\times r$) are substantially smaller ($r\ll d_{h}$), the total bytes transferred is $2L(d_{h}+Hr)$ for LRKV versus $2LHd_{h}$ for standard MHA, matching the cache-size reduction and directly translating memory savings into reduced bandwidth pressure.
Parameter complexity. Standard per-layer K/V parameters scale as $P_{\text{MHA,KV}}=2Hdd_{h}$. LRKV uses a shared base plus low-rank factors:
which may be lower or higher than standard depending on $(H,r,d_{h})$. In practice, we choose $r$ to prioritize KV-cache reduction with minimal quality impact; in our experiments, LRKV remains competitive or superior even when the total K/V parameter count is below that of standard MHA.
Rank as a control for diversity. The residual rank $r$ controls a continuous spectrum between fully shared KVs (MQA-style, $r=0$) and fully independent per-head projections (standard MHA, large $r$). The shared base $\mathbf{W}_{\mathrm{shared}}^{K,V}$ captures globally useful features, while low-rank residuals enable head specialization within a constrained budget. Empirically, we find $r\approx 0.36$-$0.43\times d_{h}$ provides a critical threshold: LRKV achieves 93.5% PCA-based effective rank versus 94.0% for standard MHA at 2.5B scale (subsection 4.3), preserving nearly all head diversity while achieving substantial cache reduction. See Appendix B for detailed discussion of spectral structure and low-rank factorization.
Figure 2: Cross-scale pretraining curves with dual-axis compute metrics. Test cross-entropy loss across four model scales (128M, 1.2B, 2.5B, 6.3B) plotted against training tokens (bottom x-axis) and cumulative training compute in ExaFLOPs (top x-axis). LRKV demonstrates competitive performance across all scales, achieving the lowest test loss at 128M and 2.5B scales. Final results in Table 1.
4 Experiments
We evaluate LRKV through large-scale pretraining experiments across four model sizes (128M, 1.2B, 2.5B, 6.3B parameters), comparing against standard MHA and state-of-the-art KV-efficient baselines. We measure pretraining loss, training efficiency, downstream task performance, and provide detailed analysis of why LRKV preserves modeling quality despite substantial cache reduction.